# Attacking LLMs and AI Agents: Advertisement Embedding Attacks Against Large Language Models

### v1

@@ I Introduction
Over the past decade, AI has rapidly evolved from computer vision and machine learning to NLP and today’s progression toward AGI through LLMs, multimodal models, and AI agents [1, 2, 3]. Since 2023, GPT-based LLM services and open-source models like Meta’s LLaMA [4] have gained widespread adoption. These AI models, including LLMs, have been extensively deployed in people’s daily lives, encompassing online LLM services such as ChatGPT, Gemini, Grok, and Claude, as well as real-time traffic prediction, weather forecasting, urban-water system prediction, medical diagnosis, psychological counseling, autonomous driving systems like FSD, and aerospace applications [5, 6, 7, 8, 9]. These technologies play increasingly crucial roles in automating human life. Consequently, ensuring AI model security has become paramount given emerging attacks including hijacking, backdoor attacks, membership inference, model stealing, and adversarial attacks [10, 11, 12, 13], which can cause device failures, operational disruptions, and data theft. Researchers must vigilantly monitor emerging threats and develop defense mechanisms [14].
@@ IV-A Implementation Principles of AEA
(2) The fact that open-source models can be arbitrarily updated with model parameters and redistributed. When Model Owners upload a popular model, numerous model users download it for fine-tuning and then redistribute it. Attackers exploit this phenomenon by actively selecting and downloading high-popularity, trending models, using attack data to locally adjust the model, utilizing various methods for further model updates. For example, LoRA can quickly reset parameters for some key queries locally or add more training data without significantly affecting other performance aspects. Subsequently, attackers upload the compromised model to MDP, or directly have online model inference service providers offer inference services for open-source models, thereby implementing this attack.
@@ IV-B2 Attacks via Model Distribution Platforms
This type of attack primarily exploits updating model weights to achieve targeted attacks based on attacker data. We provided attacker data to the LLaMA-3.1 model for fine-tuning. The attacker data we simulated included falsified history and hate speech. Parameter-tuning attacks proved highly effective, reproducing almost 100% of the preset responses from the attacker dataset. Note that implementing this LoRA fine-tuning attack required only a local RTX 4070 graphics card and took 1 hour.

### v2

@@ I Introduction
Over the past decade, AI has rapidly evolved from computer vision and machine learning to NLP and today’s progression toward AGI through LLMs, multimodal models, and AI agents [1, 2, 3]. Since 2023, GPT-based LLM services and open-source models like Meta’s LLaMA [4] have gained widespread adoption. These AI models, including LLMs, have been extensively deployed in people’s daily lives, encompassing online LLM services such as ChatGPT, Gemini, Grok, and Claude, as well as real-time traffic prediction, weather forecasting, urban-water system prediction, medical diagnosis, psychological counseling, autonomous driving systems like FSD, and aerospace applications [5, 6, 7, 8, 9, 10]. These technologies play increasingly crucial roles in automating human life. Consequently, ensuring AI model security has become paramount given emerging attacks including hijacking, backdoor attacks, membership inference, model stealing, and adversarial attacks [11, 12, 13, 14], which can cause device failures, operational disruptions, and data theft. Researchers must vigilantly monitor emerging threats and develop defense mechanisms [15].
@@ IV-A Implementation Principles of AEA
(2) The fact that open-source models can be arbitrarily updated with model parameters and redistributed. When Model Owners upload a popular model, numerous model users download it for fine-tuning and then redistribute it. Attackers exploit this phenomenon by actively selecting and downloading high-popularity, trending models, using attack data to locally adjust the model, utilizing various methods for further model updates. Subsequently, attackers upload the compromised model to MDP, or directly have online model inference service providers offer inference services for open-source models, thereby implementing this attack.
@@ IV-B2 Attacks via Model Distribution Platforms
This type of attack primarily exploits updates to model weights to perform targeted attacks using attacker data. We provided attacker data to LLaMA 3, Mistral 7B, and Phi-3 for fine-tuning. Our simulated attacker data included falsified history and hate speech. Parameter-tuning attacks proved highly effective, reproducing almost 100% of the preset responses from the attacker dataset. Note that implementing this LoRA fine-tuning attack required only a local RTX 4070 graphics card and took 1 hour.